Please answer this question it`s urgent Sinx + sin3x + sin5x = 0

Grade:11

1 Answers

Laltesh Singh
6 years ago
Sinx + sin3x + sin5x = 0Sin5x + sinx + sin3x = 02 sin3x cos2x + sin3x = 0sin3x ( 2 cosx + 1) = 0sin3x = 0 or cos2x = -1/23x = nπ or 2x = 2nπ +-2π/3 x = nπ/3 or x = nπ +- π/3